# TokenRec: Learning to Tokenize ID for LLM-based Generative Recommendations

### v1

@@ 3.3. Generalizability Evaluation
For this analysis, we consider the Beauty and LastFM datasets and exclude the 5% of users (ref as unseen users) with the least interaction history from the training data split to simulate newly added users. In this evaluation, the performance of TokenRec is compared to other LLM-based recommendation models, i.e., P5, POD, CID, and CoLLM. To ensure the absence of data leakage concerning unseen users, only the training split is utilized for establishing these LLM4Rec models, while it is allowed to update the vector database to provide collaborative representations of unseen users and items (as shown in Figure 3), which consumes far less computational resources than updating the LLM backbone of LLM-based RecSys. From the recommendation results presented in Table 4, several key observations can be made:
@@
Table 4. Performance comparison on seen and unseen users for generalizability evaluation.
@@
3.4. Efficiency Evaluation
In this subsection, we analyze the inference efficiency of TokenRec compared with the general LLM-based recommendation methods (i.e., P5, CID, and POD), which generate the tokens (e.g., IDs, titles, and descriptions) for Top-$K$ items by using auto-regressive decoding and beam search. Our proposed method discards this text-decoding generation solution and introduces a generative retrieval paradigm to perform collaborative recommendations. The results, as presented in Table 5, demonstrate that TokenRec can achieve superior inference efficiency with a significant improvement of approximately 1306.5% compared to the LLM-based recommendation baselines. This can be attributed to our generative retrieval paradigm, bypassing the most time-consuming auto-regressive decoding and beam search processes of LLMs (Wang et al., 2024b).
Table 5. Average inference time (milliseconds) per user.
@@ 3.4. Efficiency Evaluation
  • * The average improvement in inference time compared to the baselines.
3.5. Ablation Study
@@
  • w/o High-Order Collaborative Knowledge (HOCK): Use Matrix Factorization (MF) to learn the collaborative representations of users and items for ID tokenization in MQ-Tokenizers. Compared to advanced GNN-based methods (e.g., LightGCN), it is challenging to capture high-order collaborative signals among user-item interactions explicitly.
@@
  • w/o Masking: Deactivate the masking operation in users&items MQ-Tokenizers.
From the ablation results in Table 6, we can have the following observations. First, each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Second, the masking operation and the $K$-way framework introduced in our approach not only enhance the generalizability, as demonstrated in Table 4, but also result in moderate performance improvements during the ablation experiment. Finally, the removal of the high-order collaborative knowledge learned by advanced GNNs resulted in a significant drop in performance. This indicates the importance of incorporating such knowledge for the alignment of LLMs and personalized recommendations.
Table 6. Results of Ablation Studies
@@
3.6. Hyper-parameter Analysis
@@
3.6.1. Effect of Masking Ratio $\rho$
@@
Figure 4. The effect of masking ratio $\rho$ under HR@20 and NDCG@20 metrics.
@@
Figure 5. The effect of the number of sub-codebooks $K$ and the number of tokens in each sub-codebook $J$ under HR@20 and NDCG@20 metrics.
3.6.2. Effect of Codebook Settings $K$ and $L$
To study whether our proposed method TokenRec can benefit from stacking more sub-codebooks and introducing more codebook tokens, we vary the numbers of the hyper-parameters $K$ and $L$ in the range of {1, 2, 3, 4, 5} and {128, 256, 512, 1024}, respectively, and report the performance on all datasets in Figure 5, with the optimal cases highlighted by black box. From the figure, we make the following observations.
@@ 4.1. Collaborative Filtering
In order to provide personalized recommendations that accord with user preferences, collaborative filtering (CF) serves as a representative technique for modeling collaborative information in recommender systems, such as user-item interactions, to capture similar user patterns and predict future interactions (Fan et al., 2021; Fan et al., 2019a). As a representative example, matrix factorization (MF) (Koren et al., 2009; Fan et al., 2018) vectorizes users and items into dense representations and models their interactions (i.e., user-item ID matrix) by calculating the inner products between the vectorized representations. Subsequently, NeuCF (He et al., 2017) incorporates neural networks with MF to decompose the use-item interactions into two low-rank matrices representing user&item embeddings. Later on, DSCF (Fan et al., 2019c) takes advantage of deep language models to enhance user representations for collaborative recommendations by capturing auxiliary information from neighbors in social networks. Due to the superior representation learning capability in graphs, Graph Neural Networks (GNNs), such as LightGCN (He et al., 2020), GTN (Fan et al., 2022), and LTGNN (Zhang et al., 2024), are proposed to capture high-order collaborative knowledge on user-item interaction graphs for enhancing the performance of recommender systems. More specifically, GNNs take advantage of the graph-structured nature of user-item interactions and model similar user behavior patterns toward items through information propagation on user-item interaction graphs. For example, GraphRec (Fan et al., 2019b, 2020) introduces a graph attention network-based framework to encode user-item interactions and user-user social relations for social recommendations. LightGCN (He et al., 2020) is introduced to largely simplify the GNN-based recommendation methods by removing feature transformation and nonlinear activation, achieving state-of-the-art prediction performance for recommendations. In addition, GTN (Fan et al., 2020) and LTGNN (Zhang et al., 2024) provide an improvement by capturing the adaptive reliability and high-order linear-time patterns of interactions, respectively.
4.2. LLM-based Recommender Systems
@@
With the rapid development of Large Language Models (LLMs), such as ChatGPT and GPT-4, notable milestones have been showcased for revolutionizing natural language processing techniques (Ding et al., 2024a; Ding et al., 2024b). Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus (Fan et al., 2024; Li et al., 2024). In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation ability, along with remarkable generalization capability and reasoning skills that facilitate LLMs to better generalize to unseen tasks and domains (Chen et al., 2024). Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems (Zhao et al., 2024; Wang et al., 2024b). Notably, to harness the distinctive capabilities of LLMs for advancing recommender systems, existing studies have actively investigated various paradigms (Zhao et al., 2023), including pre-training, fine-tuning, and prompting (e.g., in-context learning, chain-of-thought, and instruction tuning), for adapting LLMs to recommendation tasks. For example, P5 (Geng et al., 2022) introduces an LLM-based recommendation model that unifies diverse recommendation tasks by multi-task prompt-based pre-training, which achieves impressive zero-shot generalization capability to unseen recommendation tasks with the help of personalized prompts.
@@
Despite effectiveness, most LLM-based recommendation methods still have an intrinsic limitation on indexing users and items’ IDs in language models. The naive approach known as Independent Indexing (IID) directly assigns special tokens to represent users and items within language models. Moreover, as a natural solution, textual title indexing is proposed to utilize textual contents (e.g., titles and description) to tokenize items using LLMs’ in-vocabulary tokens (Gao et al., 2023; Bao et al., 2023), such as the example ”Apple iPhone 15, 256 GB, black”. P5 (Geng et al., 2022) and POD (Li et al., 2023) apply positional and whole-word embeddings to highlight the tokens representing items and users. Most recently, several studies borrow the conception of soft prompt and utilize exogenous tokens with continuous embeddings to represent users and items in LLM-based recommendations (Zhang et al., 2023; Liao et al., 2023). Although the use of title indexing and whole-word embeddings can mitigate the issue of vocabulary explosion, it fails to effectively capture high-order collaborative knowledge and lacks generalizability for recommending to unseen users or items, as seen in P5 (Geng et al., 2022) and POD (Li et al., 2023). Furthermore, the discrete nature of tokens in language models presents a challenge in achieving tight alignment of LLMs in recommendations when utilizing continuous representations. Additionally, most LLM-based RecSys suffer from time-consuming inference due to auto-regressive decoding and beam search. However, the majority of existing methods for users&items tokenization fail to effectively capture high-order collaborative knowledge in personalized recommendations. Thus, this paper proposes a novel framework for LLM-based recommendations, TokenRec, which not only introduces a generalizable ID tokenization strategy to capture high-order collaborative knowledge but also proposes a generative retrieval paradigm to generate top-$K$ items efficiently.
5. Conclusion
@@
While existing large language model based recommendation methods achieve promising prediction performance, they fail to capture high-order collaborative knowledge and suffer from inferior generalization capability for tokenizing users&items. Additionally, the time-consuming inference remains as the emerging challenge in LLM-based recommender systems. To tackle these shortcomings, we propose a novel approach, named TokenRec, which not only introduces a generalizable user&item ID tokenization strategy to capture high-order collaborative knowledge but also presents a generative retrieval paradigm for the efficient generation of top-$K$ items. Particularly, Masked Vector-Quantized Tokenizer (MQ-Tokenizer) is developed to tokenize users and items in LLM-based recommendations by incorporating high-order collaborative knowledge. Through comprehensive experiments on four distinct datasets, we demonstrated that our model can achieve state-of-the-art recommendation performance while also exhibiting the capacity to generalize to unseen users.

### v2

@@ III-C Generalizability Evaluation
For this analysis, we consider the Beauty and LastFM datasets and exclude the 5% of users (ref as unseen users) with the least interaction history from the training data split to simulate newly added users. In this evaluation, the performance of TokenRec is compared to other LLM-based recommendation models, i.e., P5, POD, CID, and CoLLM. To ensure the absence of data leakage concerning unseen users, only the training split is utilized for establishing these LLM4Rec models, while it is allowed to update the vector database to provide collaborative representations of unseen users and items (as shown in Figure 3), which consumes far less computational resources than updating the LLM backbone of LLM-based RecSys. From the recommendation results presented in Table IV, several key observations can be made:
@@
TABLE IV: Performance comparison on seen and unseen users for generalizability evaluation.
@@
III-D Efficiency Evaluation
In this subsection, we analyze the inference efficiency of TokenRec compared with the general LLM-based recommendation methods (i.e., P5, CID, and POD), which generate the tokens (e.g., IDs, titles, and descriptions) for Top-$K$ items by using auto-regressive decoding and beam search. Our proposed method discards this text-decoding generation solution and introduces a generative retrieval paradigm to perform collaborative recommendations. The results, as presented in Table V, demonstrate that TokenRec can achieve superior inference efficiency with a significant improvement of approximately 1306.5% compared to the LLM-based recommendation baselines. This can be attributed to our generative retrieval paradigm, bypassing the most time-consuming auto-regressive decoding and beam search processes of LLMs [25].
TABLE V: Average inference time (milliseconds) per user.
@@ III-D Efficiency Evaluation
  • * The average improvement compared to the baselines.
III-E Ablation Study
@@
  • w/o High-Order Collaborative Knowledge (HOCK): Use Matrix Factorization (MF) to learn the collaborative representations of users and items for ID tokenization in MQ-Tokenizers. Compared to advanced GNN-based methods (e.g., LightGCN), it is challenging for MF to capture high-order collaborative signals explicitly among user-item interactions.
@@
  • w/o Masking: Deactivate the masking operation in our user and item MQ-Tokenizers.
We can have the following observations from the ablation results in Table VI. First, each component in our approach contributes to the overall performance since eliminating any of them would result in performance degradation. Second, the masking operation and the $K$-way framework introduced in our approach not only enhance the generalizability, as demonstrated in Table IV, but also result in moderate performance improvements during the ablation experiment. Finally, the removal of the high-order collaborative knowledge learned by advanced GNNs resulted in a significant drop in performance. This indicates the importance of incorporating such knowledge for the alignment of LLMs and personalized recommendations.
TABLE VI: Results of Ablation Studies
@@
III-F Hyper-parameter Analysis
@@
III-F1 Effect of Masking Ratio $\rho$
@@
Figure 4: The effect of masking ratio $\rho$ under HR@20 and NDCG@20 metrics.
@@
Figure 5: The effect of the number of sub-codebooks $K$ and the number of tokens in each sub-codebook $J$ under HR@20 and NDCG@20 metrics.
III-F2 Effect of Codebook Settings $K$ and $L$
To study whether our proposed method TokenRec can benefit from stacking more sub-codebooks and introducing more codebook tokens, we vary the numbers of the hyper-parameters $K$ and $L$ in the range of {1, 2, 3, 4, 5} and {128, 256, 512, 1024}, respectively, and report the performance on all datasets in Figure 5, with the optimal cases highlighted by black box. From the figure, we can make the following observations.
@@ IV-A Collaborative Filtering
In order to provide personalized recommendations that accord with user preferences, collaborative filtering (CF) serves as a representative technique for modeling collaborative information in recommender systems, such as user-item interactions, to capture similar user patterns and predict future interactions [41, 42]. As a representative example, matrix factorization (MF) [4, 43] vectorizes users and items into dense representations and models their interactions (i.e., user-item ID matrix) by calculating the inner products between the vectorized representations. Subsequently, NeuCF [34] incorporates neural networks with MF to decompose the use-item interactions into two low-rank matrices representing user&item embeddings. Later on, DSCF [44] takes advantage of deep language models to enhance user representations for collaborative recommendations by capturing auxiliary information from neighbors in social networks. Due to the superior representation learning capability in graphs, Graph Neural Networks (GNNs), such as LightGCN [5], GTN [6], and LTGNN [29], are proposed to capture high-order collaborative knowledge on user-item interaction graphs for enhancing the performance of recommender systems. More specifically, GNNs take advantage of the graph-structured nature of user-item interactions and model similar user behavior patterns toward items through information propagation on user-item interaction graphs. For example, GraphRec [3, 45] introduces a graph attention network-based framework to encode user-item interactions and user-user social relations for social recommendations. LightGCN [5] is introduced to largely simplify the GNN-based recommendation methods by removing feature transformation and nonlinear activation, achieving state-of-the-art prediction performance for recommendations. In addition, GTN [45] and LTGNN [29] provide an improvement by capturing the adaptive reliability and high-order linear-time patterns of interactions, respectively.
IV-B LLM-based Recommender Systems
@@
With the rapid development of Large Language Models (LLMs), such as ChatGPT and GPT-4, notable milestones have been showcased for revolutionizing natural language processing techniques [46, 47]. Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus [48, 49]. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation ability, along with remarkable generalization capability and reasoning skills that facilitate LLMs to better generalize to unseen tasks and domains [50]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [2, 25]. Notably, to harness the distinctive capabilities of LLMs for advancing recommender systems, existing studies have actively investigated various paradigms [8], including pre-training, fine-tuning, and prompting (e.g., in-context learning, chain-of-thought, and instruction tuning), for adapting LLMs to recommendation tasks. For example, P5 [10] introduces an LLM-based recommendation model that unifies diverse recommendation tasks by multi-task prompt-based pre-training, which achieves impressive zero-shot generalization capability to unseen recommendation tasks with the help of personalized prompts.
@@
Despite effectiveness, most LLM-based recommendation methods still have an intrinsic limitation on indexing users and items’ IDs in language models. The naive approach known as Independent Indexing (IID) directly assigns special tokens to represent users and items within language models. Moreover, as a natural solution, textual title indexing is proposed to utilize textual contents (e.g., titles and description) to tokenize items using LLMs’ in-vocabulary tokens [9, 11], such as the example ”Apple iPhone 15, 256 GB, black”. P5 [10] and POD [39] apply positional and whole-word embeddings to highlight the tokens representing items and users. Most recently, several studies borrow the conception of soft prompt and utilize exogenous tokens with continuous embeddings to represent users and items in LLM-based recommendations [14, 13]. Although the use of title indexing and whole-word embeddings can mitigate the issue of vocabulary explosion, it fails to capture high-order collaborative knowledge effectively and lacks generalizability for recommending to unseen users or items, as seen in P5 [10] and POD [39]. Furthermore, the discrete nature of tokens in language models presents a challenge in achieving tight alignment of LLMs in recommendations when utilizing continuous representations. Additionally, most LLM-based RecSys suffer from time-consuming inference due to auto-regressive decoding and beam search. However, the majority of existing methods for users&items tokenization fail to capture high-order collaborative knowledge in personalized recommendations effectively. Thus, this paper proposes a novel framework for LLM-based recommendations, TokenRec, which not only introduces a generalizable ID tokenization strategy to capture high-order collaborative knowledge but also proposes a generative retrieval paradigm to generate top-$K$ items efficiently.
V Conclusion
@@
While existing large language model-based recommendation methods achieve promising prediction performance, they fail to capture high-order collaborative knowledge and suffer from inferior generalization capability for tokenizing users&items. Additionally, the time-consuming inference remains an emerging challenge in LLM-based recommender systems. To tackle these shortcomings, we propose a novel approach, named TokenRec, which not only introduces a generalizable user&item ID tokenization strategy to capture high-order collaborative knowledge but also presents a generative retrieval paradigm for the efficient generation of top-$K$ items. Particularly, a Masked Vector-Quantized Tokenizer (MQ-Tokenizer) is developed to tokenize users and items in LLM-based recommendations by incorporating high-order collaborative knowledge. Through comprehensive experiments on four distinct datasets, we demonstrated that our model can achieve state-of-the-art recommendation performance while also exhibiting the capacity to generalize to unseen users.